# FloodDiffusion: Tailored Diffusion Forcing for Streaming Motion Generation

### v1

@@ 1 Introduction
Streaming motion generation from text prompts has growing interest due to its potential for controlling real-time agents, such as in-game NPCs or robotic systems [kodaira2023streamdiffusion, liu2022disco, yang2025shallow, liu2024tango, liu2025video, jiang2023motiongpt]. However, most existing works focus on non-streaming motion generation (often noted as text-to-motion [tevet2022human, guo2022tm2t, zhang2022motiondiffuse, petrovich2022temos, ren2022diffusionmotion]), which generates a complete motion sequence from a static text prompt. Directly concatenating outputs from such models for streaming purposes is non-viable, as it results in unsmooth transitions and significant latency.
More recent works have begun to explore dedicated streaming frameworks, primarily based on chunk-by-chunk diffusion [zhang2025primal, xie2025progressive] or auto-regressive (AR) models with a diffusion head [xiao2025motionstreamer, yin2025fastautoregressive]. For instance, PRIMAL [zhang2025primal] generates sequences via chunk-by-chunk diffusion, conditioned on past motion and the current text. MotionStreamer [xiao2025motionstreamer] uses an AR structure to handle long-term relationships, generating a single token at each step via a small diffusion head. However, these approaches have limitations: (i) chunk-by-chunk diffusion suffers from high “first-token” latency, as the generation process must wait to fill a full context length; and (ii) AR-based frameworks are limited in their ability to capture and utilize explicit history from past motions.
In the context of temporal sequence generation, another framework, diffusion forcing, was originally proposed for video generation [ho2022videodiffusion, yang2022diffusionvideogeneration, yang2023localglobal, xie2025progressive, yin2025fastautoregressive]. By assigning different time-steps to each noisy frame, it theoretically offers advantages in both “first-token” latency and the direct utilization of explicit history frames. These advantages directly address the primary limitations of existing methods. Thus, in this paper, we explore a diffusion forcing approach for the streaming human motion generation task.
@@ 2.1 Streaming Generation with Diffusion Forcing
Diffusion forcing was first described as a way to let different frames/tokens carry different noise levels so that generation can proceed in a flexible manner [benoit2025diffusionforcing]. Follow-up works reduce the train–test gap by explicitly rolling out the model during training, e.g., self forcing [huang2025selfforcing], and by keeping a longer denoising buffer [liu2025rollingforcing]. Recent works [sun2025ardiffusion, chen2025skyreel] impose a non-decreasing timestep constraint to enforce causal ordering, while still using random schedule. There are also system-level pipelines for real-time interactive diffusion that focus on latency and cache reuse [kodaira2023streamdiffusion]. These studies are mostly in video form, assume spatially large inputs, and use attention masks designed for video. Our task is different: motion is 1D in time, the control (text) may change at any step, and the model must immediately reflect the newest text on the buffered frames. Directly using the video-style diffusion forcing leads to suboptimal motion quality.
@@ 2.2 Human Motion Generation
Text-to-motion has been mainly studied in the non-streaming setting. Early works show that motion and text can be modeled jointly, either by tokenizing motion or by learning a shared latent space [guo2022tm2t, petrovich2022temos]. Diffusion-based motion models (e.g., MDM, MotionDiffuse) bring in the diffusion based methods on HumanML3D and KIT-ML [tevet2022human, zhang2022motiondiffuse, ren2022diffusionmotion]. Large or unified models further integrate text-to-motion, motion-to-text, and editing [jiang2023motiongpt, chi2024m2d2m], and discrete/token-based designs like T2M-GPT and MoMask improve compactness and FID [zhang2023t2mgpt, guo2024momask, zhang2023remodiffuse, chen2024mmdm]. These methods are usually trained and evaluated on HumanML3D and KIT-ML [guo2022generating, plappert2016kit] and assume the whole text prompt of future motion is available. More recent works start to consider online/streaming motion. MotionStreamer builds a causal latent space and generates motion step by step [xiao2025motionstreamer]. PRIMAL adopts an interactive, avatar-oriented formulation and repeatedly uses a chunk diffusion to extend motion [zhang2025primal]. They can run online, but they either suffer from first-token latency (need to fill a chunk) or only implicitly use long motion history. Our work show that a diffusion-forcing style model, once tailored for motion, can reach the quality of non-streaming text-to-motion models while staying streamable.
@@ 3.2 Vectorized Time Schedule
This schedule creates a cascading activation pattern where each frame progressively transitions from noise to data, enabling streaming generation (illustrated in Figure X).
@@ 3.4 Network Architecture
Our tailored diffusion forcing is instantiated on top of a latent diffusion formulation [rombach2022high], where we first train a causal variational autoencoder (VAE) to map motion sequences into a compact latent space and then run diffusion only in this latent space. This keeps the streaming latency low and lets the denoiser focus on temporal structure.
@@
Different from non-streaming motion works that use VAE or VQ-VAE with bi-directional convolutions [kingma2014auto, van2017neural, guo2024momask], we adopt a strictly causal design, similar in spirit to the streaming encoder in MotionStreamer [xiao2025motionstreamer]: the decoder at time $t$ does not rely on future frames. Concretely, we start from the causal VAE design used in Wan2.1 for video generation [wan2025open] and adapt all spatio-temporal blocks to 1D temporal motion sequences. Compared to Wan2.1, our version is simplified: we train it with an $\ell_{2}$ reconstruction loss and a standard commitment/codebook loss, so that the encoder produces stable latents for the downstream diffusion task.
@@
For the latent denoiser, we build on the diffusion transformer (DiT) style backbone [peebles2023scalable] as adopted in Wan2.1 [wan2025open], i.e., a single time-embedding pathway shared across blocks, instead of per-block time MLPs. Unlike the original setting for video generation, we use uniform timestep sampling and set the flow-matching time shift to $1$ during training to match our lower-triangular forcing schedule. Text conditioning is applied per frame: we extract token features from a pretrained T5 encoder (max length 128) [raffel2020exploring], flatten them to 1D, and apply the same rotary embedding to align them with the motion tokens at the current time. Inside attention, each motion frame is only allowed to attend to the text prompt active at that time step via a biased mask. This keeps self-attention over motion clean, while still injecting the latest text, and empirically improves temporal smoothness and text–motion alignment. We provide our pipeline overview in Fig. 2.
@@ 4 Experiments
We evaluate on the two standard text-conditioned motion datasets: HumanML3D [guo2022generating] and BABEL [punnakkal2021babel]. HumanML3D is built on AMASS and HumanAct12 motions and contains about 14k text–motion pairs (roughly 26 hours), each clip paired with 1–3 natural-language descriptions. We use the provided $20$ fps and joint configuration. As in prior work [tevet2022human, guo2024momask], we adopt the 263-dimensional motion representation distributed with HumanML3D (global velocity, root rotation, joint rotations, foot contact), and use the released text/motion encoders for metrics. BABEL [punnakkal2021babel] is an action-centric annotation of AMASS motion, where long motion sequences are segmented and annotated with frame-level action labels (e.g., walking $\rightarrow$ turn $\rightarrow$ jump). we follow the official train/val/test split released with these dataset. Concretely, we follow the official BABEL split and extract continuous multi-action segments to form text–motion pairs, similar to [xiao2025motionstreamer]. This setup matches our streaming assumption, i.e., the text prompt may change inside a single motion sequence.
@@
We follow the evaluation protocol widely used in text-to-motion [guo2022generating, tevet2022human, guo2024momask]. We evaluate Fréchet Inception Distance (FID), R-Precision, Diversity, and Multimodality Distance (MM-Dist), Peak Jerk (PJ) and Area Under the Jerk (AUJ). For FID, we adopt the same motion encoderas HumanML3D [guo2022generating] for encoding the real motions and generated motions. For R-Precision (top-$k$), given a generated motion, we encode it and retrieve among ground-truth text description plus mismatched texts. If the ground-truth text is within the top-$k$ ranked texts, it is counted as correct. We report R-Precision@1/2/3, following guo2022generating, tevet2022human. For Diversity, we measure whether the model collapses by randomly sampling several generated motions, and computing the average pairwise $\ell_{2}$ distance. For Multimodality Distance (MM-Dist), we measure the distance between motions and texts. For PJ [barquero2024seamless], we calculate the maximum value throughout the transition motion over all joints, and similarly, AUJ[barquero2024seamless] is for the area under the jerk curve.
@@
TM2T [guo2022tm2t] 0.424±.003 0.618±.003 0.729±.002 1.501±.017 3.467±.011 8.589±.076 - -
T2M [guo2022generating] 0.455±.003 0.636±.003 0.736±.002 1.087±.021 3.347±.008 9.175±.083 - -
MDM [tevet2022human] - - 0.611±.007 0.544±.044 5.566±.027 9.559±.086 - -
MLD [chen2023executing] 0.481±.003 0.673±.003 0.772±.002 0.473±.013 3.196±.010 9.724±.082 - -
MotionDiffuse [zhang2022motiondiffuse] 0.491±.001 0.681±.001 0.782±.001 0.630±.001 3.113±.001 9.410±.049 - -
T2M-GPT [zhang2023t2mgpt] 0.492±.003 0.679±.002 0.775±.002 0.141±.005 3.121±.009 9.722±.082 -
ReMoDiffuse [zhang2023remodiffuse] 0.510±.005 0.698±.006 0.795±.004 0.103±.004 2.974±.016 9.018±.075 - -
MoMask [guo2024momask] 0.521±.002 0.713±.002 0.807±.002 0.045±.002 2.958±.008 9.677±.032 - -
PRIMAL [zhang2025primal] 0.497±.003 0.681±.003 0.780±.002 0.511±.006 3.120±.019 9.520±.068 1.304±.105 19.36±.969
MotionStreamer [xiao2025motionstreamer] 0.513±.003 0.705±.006 0.802±.003 0.092±.003 2.909±.015 9.722±.037 0.912±.049 16.57±.762
@@
MotionStreamer [xiao2025motionstreamer] -0.338 -0.136 -0.055
PRIMAL [zhang2025primal] -0.599 -0.315 -0.203
@@
VQ-VAE [zhang2023t2mgpt] 0.0062 0.0545 0.0341 0.1259
CausalVAE [xiao2025motionstreamer] 0.0103 0.0576 0.0351 0.0271
Wan CausalVAE [wan2025open] 0.0094 0.0555 0.0333 0.0290
Table 4: Ablation on Causal VAE architectures. We compare VAE reconstruction quality using VQ-VAE [zhang2023t2mgpt], MotionStreamer’s CausalVAE [xiao2025motionstreamer], and our adopted Wan CausalVAE [wan2025open]. The results show comparable performance across the different causal variants.
@@
We first train on HumanML3D alone to evaluate the standard text-to-motion setting, where every frame in a clip shares the same text description, following [guo2022generating, tevet2022human]. Then we train a joint model on HumanML3D and BABEL [punnakkal2021babel] to evaluate scenarios with time-varying prompts; in this setting, each frame (or short segment) is conditioned on its corresponding text, requiring the model to update the motion immediately as the text changes. In both training settings, the model operates in streaming mode, denoising the current buffer and enabling frame-by-frame output without waiting for the full sequence.
@@ 4.1 Comparison to state-of-the-art approaches
We first train our model on HumanML3D for non-streaming comparison, where the input is a single text and the output is a complete motion matching the ground-truth length. We compare our method with previous state-of-the-art models up to MoMask[guo2024momask], including the retrieval-based method ReMoDiffuse[zhang2023remodiffuse], the auto-regressive (AR) method T2M-GPT [zhang2023t2mgpt], and the diffusion-based method MDM [tevet2022human]. Then, we train our model on jointly on HumanML3D and the BABEL dataset from scratch for streaming comparison. Following previous works [zhang2025primal, xiao2025motionstreamer], the evaluation inputs past motion, past text, and current text to generate the subsequent motion. We reproduce PRIMAL [zhang2025primal] and MotionStreamer [xiao2025motionstreamer] using their official code, adapting them to the 263D [guo2022generating] HumanML3D motion representation to ensure a consistent comparison with our non-streaming results. We note that all frameworks could be trained on other motion representations, such as the 272D representation in MotionStreamer [xiao2025motionstreamer] or the 330D SMPLX [SMPL-X:2019] 6D-rotation [zhou2019continuity] representation.
@@
As shown in Table 1, our method performs better on most non-streaming metrics and leads on all streaming metrics. For the non-streaming evaluation, our method outperforms existing baselines in R-Precision and Multimodal Distance, and performs on par with MoMask[guo2024momask] in FID. For streaming-specific metrics, our method achieves a PJ closest to real motion and performs better on AUJ. Overall, the results demonstrate our method can generate streaming motion at a state-of-the-art quality.
@@ 4.2 Ablation Study
We first evaluate the importance of bi-directional attention, as shown in Table 3. Unlike full-chunk diffusion models [zhang2025primal], where switching from bi-directional to causal attention partially influences the results (e.g., FID increasing from 0.51 to 0.92), we found bi-directional attention is critical for diffusion forcing. In our framework, using causal attention causes the FID to degrade significantly, from 0.054 to 3.377. This suggests that the denoising process requires the full context of the active window, especially when frames within that window are at different timesteps.
@@ Appendix E Details of User Study
To evaluate the perceptual quality of the generated motions, we conducted a user study with 100 participants. We compared our FloodDiffusion against real motion ground truth (GT) and two streaming baselines: PRIMAL [zhang2025primal] and MotionStreamer [xiao2025motionstreamer].

### v2

@@ 1 Introduction
Streaming motion generation from text prompts has growing interest due to its potential for controlling real-time agents, such as in-game NPCs or robotic systems [13, 14, 32, 16, 15, 11]. However, most existing works focus on non-streaming motion generation (often noted as text-to-motion [27, 8, 36, 20, 24]), which generates a complete motion sequence from a static text prompt. Directly concatenating outputs from such models for streaming purposes is non-viable, as it results in unsmooth transitions and significant latency.
More recent works have begun to explore dedicated streaming frameworks, primarily based on chunk-by-chunk diffusion [38, 31] or auto-regressive (AR) models with a diffusion head [30, 34]. For instance, PRIMAL [38] generates sequences via chunk-by-chunk diffusion, conditioned on past motion and the current text. MotionStreamer [30] uses an AR structure to handle long-term relationships, generating a single token at each step via a small diffusion head. However, these approaches have limitations: (i) chunk-by-chunk diffusion suffers from high “first-token” latency, as the generation process must wait to fill a full context length; and (ii) AR-based frameworks are limited in their ability to capture and utilize explicit history from past motions.
In the context of temporal sequence generation, another framework, diffusion forcing, was originally proposed for video generation [9, 33, 31, 34]. By assigning different time-steps to each noisy frame, it theoretically offers advantages in both “first-token” latency and the direct utilization of explicit history frames. These advantages directly address the primary limitations of existing methods. Thus, in this paper, we explore a diffusion forcing approach for the streaming human motion generation task.
@@ 2.1 Streaming Generation with Diffusion Forcing
Diffusion forcing was first described as a way to let different frames/tokens carry different noise levels so that generation can proceed in a flexible manner [2]. Follow-up works reduce the train–test gap by explicitly rolling out the model during training, e.g., self forcing [10], and by keeping a longer denoising buffer [17]. Recent works [26, 3] impose a non-decreasing timestep constraint to enforce causal ordering, while still using random schedule. There are also system-level pipelines for real-time interactive diffusion that focus on latency and cache reuse [13]. These studies are mostly in video form, assume spatially large inputs, and use attention masks designed for video. Our task is different: motion is 1D in time, the control (text) may change at any step, and the model must immediately reflect the newest text on the buffered frames. Directly using the video-style diffusion forcing leads to suboptimal motion quality.
@@ 2.2 Human Motion Generation
Text-to-motion has been mainly studied in the non-streaming setting. Early works show that motion and text can be modeled jointly, either by tokenizing motion or by learning a shared latent space [8, 20]. Diffusion-based motion models (e.g., MDM, MotionDiffuse) bring in the diffusion based methods on HumanML3D and KIT-ML [27, 36, 24]. Large or unified models further integrate text-to-motion, motion-to-text, and editing [11], and discrete/token-based designs like T2M-GPT and MoMask improve compactness and FID [35, 6, 37, 5]. These methods are usually trained and evaluated on HumanML3D and KIT-ML [7, 21] and assume the whole text prompt of future motion is available. More recent works start to consider online/streaming motion. MotionStreamer builds a causal latent space and generates motion step by step [30]. PRIMAL adopts an interactive, avatar-oriented formulation and repeatedly uses a chunk diffusion to extend motion [38]. They can run online, but they either suffer from first-token latency (need to fill a chunk) or only implicitly use long motion history. Our work show that a diffusion-forcing style model, once tailored for motion, can reach the quality of non-streaming text-to-motion models while staying streamable.
@@ 3.2 Vectorized Time Schedule
This schedule creates a cascading activation pattern where each frame progressively transitions from noise to data, enabling streaming generation (illustrated in Figure 3).
@@ 3.4 Network Architecture
Our tailored diffusion forcing is instantiated on top of a latent diffusion formulation [25], where we first train a causal variational autoencoder (VAE) to map motion sequences into a compact latent space and then run diffusion only in this latent space. This keeps the streaming latency low and lets the denoiser focus on temporal structure.
@@
Different from non-streaming motion works that use VAE or VQ-VAE with bi-directional convolutions [12, 28, 6], we adopt a strictly causal design, similar in spirit to the streaming encoder in MotionStreamer [30]: the decoder at time $t$ does not rely on future frames. Concretely, we start from the causal VAE design used in Wan2.1 for video generation [29] and adapt all spatio-temporal blocks to 1D temporal motion sequences. Compared to Wan2.1, our version is simplified: we train it with an $\ell_{2}$ reconstruction loss and a standard commitment/codebook loss, so that the encoder produces stable latents for the downstream diffusion task.
@@
For the latent denoiser, we build on the diffusion transformer (DiT) style backbone [19] as adopted in Wan2.1 [29], i.e., a single time-embedding pathway shared across blocks, instead of per-block time MLPs. Unlike the original setting for video generation, we use uniform timestep sampling and set the flow-matching time shift to $1$ during training to match our lower-triangular forcing schedule. Text conditioning is applied per frame: we extract token features from a pretrained T5 encoder (max length 128) [23], flatten them to 1D, and apply the same rotary embedding to align them with the motion tokens at the current time. Inside attention, each motion frame is only allowed to attend to the text prompt active at that time step via a biased mask. This keeps self-attention over motion clean, while still injecting the latest text, and empirically improves temporal smoothness and text–motion alignment. We provide our pipeline overview in Fig. 2.
@@ 4 Experiments
We evaluate on the two standard text-conditioned motion datasets: HumanML3D [7] and BABEL [22]. HumanML3D is built on AMASS and HumanAct12 motions and contains about 14k text–motion pairs (roughly 26 hours), each clip paired with 1–3 natural-language descriptions. We use the provided $20$ fps and joint configuration. As in prior work [27, 6], we adopt the 263-dimensional motion representation distributed with HumanML3D (global velocity, root rotation, joint rotations, foot contact), and use the released text/motion encoders for metrics. BABEL [22] is an action-centric annotation of AMASS motion, where long motion sequences are segmented and annotated with frame-level action labels (e.g., walking $\rightarrow$ turn $\rightarrow$ jump). we follow the official train/val/test split released with these dataset. Concretely, we follow the official BABEL split and extract continuous multi-action segments to form text–motion pairs, similar to [30]. This setup matches our streaming assumption, i.e., the text prompt may change inside a single motion sequence.
@@
We follow the evaluation protocol widely used in text-to-motion [7, 27, 6]. We evaluate Fréchet Inception Distance (FID), R-Precision, Diversity, and Multimodality Distance (MM-Dist), Peak Jerk (PJ) and Area Under the Jerk (AUJ). For FID, we adopt the same motion encoderas HumanML3D [7] for encoding the real motions and generated motions. For R-Precision (top-$k$), given a generated motion, we encode it and retrieve among ground-truth text description plus mismatched texts. If the ground-truth text is within the top-$k$ ranked texts, it is counted as correct. We report R-Precision@1/2/3, following Guo et al. [7], Tevet et al. [27]. For Diversity, we measure whether the model collapses by randomly sampling several generated motions, and computing the average pairwise $\ell_{2}$ distance. For Multimodality Distance (MM-Dist), we measure the distance between motions and texts. For PJ [1], we calculate the maximum value throughout the transition motion over all joints, and similarly, AUJ[1] is for the area under the jerk curve.
@@
TM2T [8] 0.424±.003 0.618±.003 0.729±.002 1.501±.017 3.467±.011 8.589±.076 - -
T2M [7] 0.455±.003 0.636±.003 0.736±.002 1.087±.021 3.347±.008 9.175±.083 - -
MDM [27] - - 0.611±.007 0.544±.044 5.566±.027 9.559±.086 - -
MLD [4] 0.481±.003 0.673±.003 0.772±.002 0.473±.013 3.196±.010 9.724±.082 - -
MotionDiffuse [36] 0.491±.001 0.681±.001 0.782±.001 0.630±.001 3.113±.001 9.410±.049 - -
T2M-GPT [35] 0.492±.003 0.679±.002 0.775±.002 0.141±.005 3.121±.009 9.722±.082 -
ReMoDiffuse [37] 0.510±.005 0.698±.006 0.795±.004 0.103±.004 2.974±.016 9.018±.075 - -
MoMask [6] 0.521±.002 0.713±.002 0.807±.002 0.045±.002 2.958±.008 9.677±.032 - -
PRIMAL [38] 0.497±.003 0.681±.003 0.780±.002 0.511±.006 3.120±.019 9.520±.068 1.304±.105 19.36±.969
MotionStreamer [30] 0.513±.003 0.705±.006 0.802±.003 0.092±.003 2.909±.015 9.722±.037 0.912±.049 16.57±.762
@@
MotionStreamer [30] -0.338 -0.136 -0.055
PRIMAL [38] -0.599 -0.315 -0.203
@@
VQ-VAE [35] 0.0062 0.0545 0.0341 0.1259
CausalVAE [30] 0.0103 0.0576 0.0351 0.0271
Wan CausalVAE [29] 0.0094 0.0555 0.0333 0.0290
Table 4: Ablation on Causal VAE architectures. We compare VAE reconstruction quality using VQ-VAE [35], MotionStreamer’s CausalVAE [30], and our adopted Wan CausalVAE [29]. The results show comparable performance across the different causal variants.
@@
We first train on HumanML3D alone to evaluate the standard text-to-motion setting, where every frame in a clip shares the same text description, following [7, 27]. Then we train a joint model on HumanML3D and BABEL [22] to evaluate scenarios with time-varying prompts; in this setting, each frame (or short segment) is conditioned on its corresponding text, requiring the model to update the motion immediately as the text changes. In both training settings, the model operates in streaming mode, denoising the current buffer and enabling frame-by-frame output without waiting for the full sequence.
@@ 4.1 Comparison to state-of-the-art approaches
We first train our model on HumanML3D for non-streaming comparison, where the input is a single text and the output is a complete motion matching the ground-truth length. We compare our method with previous state-of-the-art models up to MoMask[6], including the retrieval-based method ReMoDiffuse[37], the auto-regressive (AR) method T2M-GPT [35], and the diffusion-based method MDM [27]. Then, we train our model on jointly on HumanML3D and the BABEL dataset from scratch for streaming comparison. Following previous works [38, 30], the evaluation inputs past motion, past text, and current text to generate the subsequent motion. We reproduce PRIMAL [38] and MotionStreamer [30] using their official code, adapting them to the 263D [7] HumanML3D motion representation to ensure a consistent comparison with our non-streaming results. We note that all frameworks could be trained on other motion representations, such as the 272D representation in MotionStreamer [30] or the 330D SMPLX [18] 6D-rotation [39] representation.
@@
As shown in Table 1, our method performs better on most non-streaming metrics and leads on all streaming metrics. For the non-streaming evaluation, our method outperforms existing baselines in R-Precision and Multimodal Distance, and performs on par with MoMask[6] in FID. For streaming-specific metrics, our method achieves a PJ closest to real motion and performs better on AUJ. Overall, the results demonstrate our method can generate streaming motion at a state-of-the-art quality.
@@ 4.2 Ablation Study
We first evaluate the importance of bi-directional attention, as shown in Table 3. Unlike full-chunk diffusion models [38], where switching from bi-directional to causal attention partially influences the results (e.g., FID increasing from 0.51 to 0.92), we found bi-directional attention is critical for diffusion forcing. In our framework, using causal attention causes the FID to degrade significantly, from 0.054 to 3.377. This suggests that the denoising process requires the full context of the active window, especially when frames within that window are at different timesteps.
@@ Appendix E Details of User Study
To evaluate the perceptual quality of the generated motions, we conducted a user study with 100 participants. We compared our FloodDiffusion against real motion ground truth (GT) and two streaming baselines: PRIMAL [38] and MotionStreamer [30].